PartnerObserving Log for 2007 Apr 10/11

Observers: John Hill, Mark Wagner
Telescope Operator: David Gonzales
Instrument: LBC-Blue

Some more SDT-related observing during engineering time. Note that I'm not paying any attention to the SDT queue. I'm only using convenient SDT targets to test various aspects of the telescope. (i.e. quicklook display script, new guiding software)

The LBC-Blue baffling and vignetting should be the same as during the February and March SDT runs. The undersized L2 baffle is responsible for the vignetting around the edge of the field.

Open at 01:55 UT. T=1.5 degC D=-7degC Wind 6m/sec from SW. Sky clear.

02:05 - 25 bias frames

02:35 - Peculiar Panic of SX Mirror (acting like it was momentarily air starved?).

03:07 - 03:26 Bfastextra on the meridian. Replacement of SX Hardpoint 4 has had no detrimental effects on the collimation table zeropoint.....maybe even improved it.

03:43 - V838 Mon -- Guiding seems to be working OK in this poor and variable seeing 2-3 arcsec. Of course, you don't really need to guide in this seeing. Wind gusting to 12 m/sec.

05:00 - M81 extrafocal 2

05:10 M81_V

05:18 M81_BV -- This set has the "best" seeing at 1.4 arcsec FWHM

05:32 - M81 extrafocal 2, focus is changing rapidly because the mirror temperature is cooling at about 1 degC per hour along with the telescope steel.

05:36 - M81_BV with seeing worse.

05:52 - M81_BV again (we are using it to test the Science.py quick display script)

06:12 - PMC Left emits "PMCLogger" errors, but mirror is still working.

06:01 Move to Egami_Lockman_VLA_B_1 for some more guider tests with 360 sec exposures.

06:17 - Egami_lockman_VLA_B_1.ob -- Dither 1 guiding failed for lack of stars. Dither 2 guiding on a galaxy. Dither 3 guiding on a star in 3 arcsec seeing. Dither 4 guiding on a star in 3 arcsec seeing --- seems to be avoiding the saturated star (or else it isn't saturated in this bad seeing). Dither 5 guiding failed for lack of stars in a 16 sec exposure (seeing is poor).

06:54 - Ufastextra40

06:57 - Egami_Lockman_VLA_U_1.ob -- Dither 1 guiding failed for lack of stars after 32 sec exposure. Stopped OB.

07:01 - Egami_Lockman_XMM_U_1.ob -- Dither 1 found a star at 32 sec exposure (2-3 arcsec seeing). Dither 2 tried guiding on a galaxy and failed legitimately. Dithers 3 and 4 found stars to guide on. Seeing is about 1.8 arcsec. Dither 5 failed to find a star and was unguided.

07:40 - Ufastextra40 at M94 field -- no good pupils for this field in U, needs a dedicated extrafocal position. 07:45 - Vfastextra32 at M94 field.

07:51 - M94_NGC_4736_U.ob -- seeing variable around 2 arcsec. Found guide stars for dithers 1,2,6,7,9,10. Failed for dithers 3,4,5,8.

Interupt Blue data taking to work on Red alignment with the Electrim camera. Our daytime surveying with the Sokkia seems to be paying some dividends. We have improved the image to have only 100 pixels of coma (more than a factor of 5 improvement since the previous image).

09:15 - Rfastextra

09:21 - 5 sec R image coincident with the Red Electrim image, but Red still has too much coma to match the star pattern. Followed by 100 sec R image.

09:26 - Vfastextra with corrections followed by 4 more Vfastextras without sending corrections. Sent another correction at 09:35.

09:38 - Bfinefoc to check filter offset. A second finefoc sequence a few minutes later.

09:56 - Vfastextra at the field of N5024. Wind has calmed to 5 m/sec, and seeing has improved to ~1 arcsec.

10:00 - NGC5024Vx3 -- Three 300-sec V exposures of NGC5024 with guiding.

10:18 - Repeat same three 300-sec exposures of NGC5024 with guiding blocked at PCS (provides tech chip images w/o guiding for comparison). Besides the systematic drift, something (wind? atmosphere? telescope?) is making small jumps (0.3 arcsec) Standardsof the image on guider (10 sec) timescales.

10:40 - Vfastextra at the field of NGC6503. Focus corrections are smaller now that the world is in better temperature equilibrium.

10:44 - NGC_6503_V with guiding turned back on. Last quarter moon is up, sky is beautifully clear. Seeing in V is 1.2 arcsec at 1.2 airmasses. Dither 5 had a bright guide star.

11:24 - SA107_Vgr standard field at elevation 49 deg.

11:30 - SA017_UB standard field at elevation 48 deg.

11:38 - SA110_Vgr standard field at elevation 53.7 deg.

11:43 - SA110_UB standard field at elevation 54.3 deg.

11:50 - Skyflat Test V 1 sec (still darkish)

11:51 - SkyflatV_5 x10 (dithered)

11:56 - SkyflatVrot180_5 x20 (dithered)

12:03 - SkyflatV_5 x30 (dithered)

12:09 - SkyFlatVrot180_5 x20 (dithered)

Close 05:20 UT T=-0.8degC, D=-14.4degC Wind 3.6 m/sec from NW. Photometric.

LBC UI had many soft hangs during the night. Most were easily recovered by changing pages, but two lasted for ~1 minute. None of the hangs required drastic measures such as rebooting.

-- JohnHill - 11 Apr 2007
Topic revision: r4 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback